In the graph of the equation ax+3y=9 is perpendicular to the graph of the equation 3x+y=-4, find the value a
Answer by Alan3354(69443)

In the graph of the equation ax+3y=9 is perpendicular to the graph of the equation 3x+y=-4, find the value a
----------------
3x+y=-4 --> y = -3x - 4 so its slope, m1, = -3
The slope of lines perpendicular will have the negative inverse slope, +1/3.
ax+3y=9
Put into slope-intercept form (that means solve for y)
3y = -ax + 9
y = (-a/3)x + 3
m2 = +1/3 = -a/3
a = -1